Status Report

SOHO Daily Meeting Minutes for Friday, 7 June 2002

By SpaceRef Editor
June 7, 2002
Filed under , ,

* RHESSI target for today is AR 9973; target for tomorrow
is TBD.

Spacecraft Status: Nominal

Spacecraft Anomalies:
157/1607 MTOPTS4 flags YL for 1 format. Experimenter contacted.

Accomplished Activities: NOM HGA, VIRGO, New Tracking Star

Planned Activities: RSL, VIRGO, SWAN

Upcoming Operations: None

Ground Anomalies:
157/2138 52 minutes degraded TM. D16. RFI. Nonrecoverable. 1
minute TM lost at 2138. Nonrecoverable.
158/0345 1.4 minutes TM dropout. D46. TGC failure. Recoverable.

06/06 DOY 157 – 06/07 DOY 158

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. RHESSI support. Weekend plans TBD.

UVCS: Nominal. Coronal hole observations at the south pole and
mini-synoptic. Weekend: coronal hole observations and synoptic.

LASCO: Nominal. C2/C3 synoptics and JOP159 support.

EIT: Nominal. 195 CME watch and synoptic.

MDI: Nominal. Full disk magnetograms and dopplergrams, doing
campaign tests today.

TRACE: Nominal. Observing AR9973 for today and the weekend.

SpaceRef staff editor.